THE EFFECT OF DEUTERIUM SUBSTITUTION IN THE AMINO/IMINO GROUP ON THE DYNAMICAL MOLECULAR STRUCTURE OF PEPTIDE MOLECULES:ACETAMIDE AND N-METHYL FORMAMIDE

\underline{EIZI HIROTA}, The Graduate University for Advanced Studies, Hayama, Kanagawa 240-0193, Japan; YOSHIYUKI KAWASHIMA, TSUYOSHI USAMI, and KOICHI SETO, Department of Applied Chemistry, Kanagawa Institute of Technology, Atsugi, Kanagawa 243-0292, Japan.

Peptide molecules XCO-NYY′ are characterized by the low potential barrier V_{3} to internal rotation of CH_{3} substituted for X and/or Y. A most conspicuous example is acetamide, for which V_{3} was previously reported to be 25.043857(19)cm^{−1}. The present study intended to clarify why V_{3} is so low in peptide molecules, by examining the effect of the out-of-plane bending or inversion of the amino/imino group on the molecular structure through deuterium substitution for amino/imino hydrogens. The potential barrier V_{3} in acetamide was found to decrease by 2.509, 2.958, and 5.422cm^{−1}, when H’s at cis, trans, and both positions in the amino group were replaced by deuterium atoms, respectively. The reduction was proportional to the effective mass of the amino inversion, which was in turn ascribed to the change in electronic resonance character of the peptide linkage. The amino inversion is coupled with the CH_{3} internal rotation, producing an interaction term proportional to τsin3α (τ: amino inversion, α: methyl internal rotation). This coupling term, when the inversion is treated by second order perturbation, yields a V_{6} term, in agreement with Ilyushin etal., who derived an unusually large V_{6} term of -10.044874(73) cm^{−1}. Deuteration for the imino hydrogen reduces V_{3} by 11.40cm^{−1}for the N-methyl formamide trans form, whereas for the cis form it increases V_{3} by 8.09cm^{−1}. It is interesting that even a small perturbation like deuterium substitution causes a substantial change in electronic structure of the peptide systems.